OpenStudy (qqstory):

radical x + radical x-7 = 21/radical x-7

OpenStudy (qqstory):

\[\sqrt{x} + \sqrt{x-7} = 21\div \sqrt{x-7}\]

OpenStudy (qqstory):

this is hard..

OpenStudy (alexwee123):

no its just algebra

OpenStudy (qqstory):

help please

OpenStudy (alexwee123):